OpenStudy (lukecrayonz):

Integrals: Please explain:) Evaluate:

OpenStudy (lukecrayonz):

\[\int\limits_{}^{}\sqrt[11]{x}*dx\]

OpenStudy (isaiah.feynman):

Know this rule? \[\int\limits x^{n} = \frac{ x^{n+1} }{ n+1 } \]

OpenStudy (isaiah.feynman):

@Lukecrayonz

OpenStudy (lukecrayonz):

Yes sorry, so just x^(1/11) 1/11+1=12/11 x^(12/11)/(12/11)?

OpenStudy (isaiah.feynman):

Yes

OpenStudy (lukecrayonz):

Gotcha, thanks man. what about for this? \[\frac{ 5 }{ x }-2e^{5x}+\sqrt[]{x^{7}}\]

OpenStudy (isaiah.feynman):

What are you required to do here?

OpenStudy (lukecrayonz):

Integral

OpenStudy (isaiah.feynman):

definite or indefinite?

OpenStudy (lukecrayonz):

Indefinite

OpenStudy (isaiah.feynman):

Use this..\[\int\limits f(x)+g(x)+h(x) dx = \int\limits f(x) dx +\int\limits g(x)dx+\int\limits h(x)dx\]

OpenStudy (lukecrayonz):

So to find the integral of 5/x, it'd be what? o.O

OpenStudy (isaiah.feynman):

Can you re write 5/x in another form?

OpenStudy (lukecrayonz):

Uhhhh yeah but I forget what form :P Is it 5^(1/x)?

OpenStudy (isaiah.feynman):

Nope.